OpenStudy (anonymous):

In a fruit basket containing apples and oranges, 1/6 of the fruit are oranges. Which could be the numbers of oranges and apples in the basket?

OpenStudy (anonymous):

is this multiply choice ?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

oops..typo...multiple choice what are the answer choices

OpenStudy (anonymous):

A. 3 oranges 15 apples B. 1 orange 6 apples C. 15 oranges 3 apples D. 6 oranges and 1 apple

OpenStudy (anonymous):

1/6 of the fruit are oranges... 3 oranges + 15 apples = 18 pieces of fruit 1/6 of 18 = 1/6(18) = 18/6 = 3 oranges answer choice A

OpenStudy (anonymous):

Thanks!

OpenStudy (anonymous):

anytime :)